Important equations for making up solutions

Weight of solute (g) = formula weight of solute (g/mole) x molarity (mol/l) x final volume (L)

C1V1 = C2V2

How to make solutions

Choose the correct formula from above. Please show all the steps and be sure to clearly show what values are being using relating to the formula.

  1. Make 80 ml of a 0.01 M KMnO4 solution.
  2. Using the 0.01 M KMnO4 stock solution make 25 ml of a 2mM KMnO4 solution.
Answer #1

For making 80mL of a 0.01M KMnO4 ​​​​​​, we need to calculate the mass of KMnO4.

Formula weight of KMnO4= sum of atomic masses of k, Mn and O

(Atomic mass of K=39g, Mn=54.9, O =16g

Formula weight of solute (KMnO4)= 39+54.9+16*4=39g+54.9+64=157.9g

one mole =formula weight =157.9g

So,we can write formula weight of KMnO4 =157.9g/mol

Using first formula,calculate weight of KMnO4

Weight of KMnO4 (g) = formula weight of KMnO4(g/mol)*molarity (mol/L) *Final volume(L)

Molarity of KMnO4  =0.01M

Volume of KMnO4= 80mL

1mL =0.001L

80mL =0.001L*80mL/1mL=0.08L

Putting all these values in the above given formula

Weight of KMnO4= 157.9g/mol*0.01mol/L*0.08L=0.12632g

After round off, weight of solute is 0.126g

Molarity is 0.01M. Molarity is mole of solute dissolved per liter of the solution. So, instead of M we can use unit mol/liter. The units mol and L will get cancelled.

2) In second case, we are diluting the solution of KMnO4 ​​​​​​.

Use the formula C1V1=C2V2

Here,C1= initial concentration

C2= final concentration

V1=initial volume

V2= final volume

Given C1=0.01M

C2 =2mM =2*10-3 M (since 1mM =10-3 M)

V1=?

V2= 25mL

C1V1=C2V2

V1=C2V2/C1= 2*10-3M *25mL/0.01M=5mL

Second formula gives us the volume of KMnO4 required  to make 25mL of 2mM KMnO4 solution.